# DeepSport: A Multimodal Large Language Model for Comprehensive Sports Video Reasoning via Agentic Reinforcement Learning

### v1

@@ Abstract
Sports video understanding presents unique challenges, requiring models to perceive high-speed dynamics, comprehend complex rules, and reason over long temporal contexts. While Multimodal Large Language Models (MLLMs) have shown promise in genral domains, the current state of research in sports remains narrowly focused: existing approaches are either single-sport centric, limited to specific tasks, or rely on training-free paradigms that lack robust, learned reasoning process. To address this gap, we introduce DeepSport, the first end-to-end trained MLLM framework designed for multi-task, multi-sport video understanding. DeepSport shifts the paradigm from passive frame processing to active, iterative reasoning, empowering the model to “think with videos” by dynamically interrogating content via a specialized frame-extraction tool. To enable this, we propose a data distillation pipeline that synthesizes high-quality Chain-of-Thought (CoT) trajectories from 10 diverse data source, creating a unified resource of 78k training data. We then employ a two-stage training strategy, Supervised Fine-Tuning (SFT) followed by Reinforcement Learning (RL) with a novel gated tool-use reward, to optimize the model’s reasoning process. Extensive experiments on the testing benchmark of 6.7k questions demonstrate that DeepSport achieves state-of-the-art performance, significantly outperforming baselines of both proprietary model and open-source models. Our work establishes a new foundation for domain-specific video reasoning to address the complexities of diverse sports.
footnotetext: Equal contribution.
@@ 1 Introduction
The domain of sports analysis has long served as a fertile ground for interdisciplinary research, driving advancements in both Computer Vision (CV) and Natural Language Processing (NLP) [THOMAS20173]. Traditional CV approaches have achieved remarkable success in specific, isolated tasks, such as fine-grained action detection and recognition [shao2020finegym, giancola2018soccernet], event spotting [Xarles_2024_CVPR], highlight generation [merler2018excitement, merler2018automatic], and player and ball trajectory detection and tracking in team sports [cioppa2021cameracalibrationplayerlocalization, vandeghen2022semi, 10411638, cite-key]. Simultaneously, NLP techniques have been effectively deployed for automated match reporting [huang2020generating] and sentiment analysis [vujivcic2023approach]. Recently, the emergence of Multimodal Large Language Models (MLLMs) [openai2024gpt4o, geminiteam2024gemini] has accelerated this trend, enabling more complex applications in sports [xia2024language], such as AI-assisted refereeing [held2023vars, held2023vars] and commentary generation [rao2024matchtimeautomaticsoccergame, you2025timesoccer].
In response to these capabilities, the community has produced a surge of high-quality, public benchmarks covering a wide array of disciplines. These range from action quality assessment in diving [xu2022finediving] to tactical classification in fencing [lai2024facts], and comprehensive question answering in large-scale benchmarks [xia2024sportqa, xia2024sportu, rao2025multi]. These datasets provide the necessary foundation for evaluating sophisticated sports intelligence.
However, despite this proliferation of data and benchmarks, the development of capable models has lagged behind. The current sports-specific MLLMs exhibits a significant fragmentation. For example, trained models are predominantly soccer-centric [rao2025multi, you2025timesoccer], optimizing for specific tasks like commentary generation within a single sport, but failing to generalize to others. On the other hand, attempts at multi-sport reasoning, such as FineQuest [chen2025finequestadaptiveknowledgeassistedsports], rely on training-free paradigms. Consequently, to the best of our knowledge, a unified, end-to-end trained MLLM capable of performing multiple tasks (e.g., commentary, foul detection, QA) across diverse sports does not yet exist.
@@
To bridge this gap, we introduce DeepSport, the first sport-specific MLLM framework explicitly trained for multi-task, multi-sport video understanding (Figure 1). Inspired by the emerging paradigm of “Thinking with Videos” [zhang2025thinkingvideosmultimodaltoolaugmented, he2025framethinkerlearningthinklong, ge2025framemindframeinterleavedvideoreasoning], DeepSport shifts from passive frame processing to active, iterative video interrogation via a specialized frame-extraction tool. To enable this, we devise a data distillation pipeline that synthesizes high-quality Chain-of-Thought (CoT) trajectories from 10 diverse data sources (derived from 9 existing works) covering 12 different sports, establishing a unified resource for training. We then employ a two-stage strategy—Supervised Fine-Tuning (SFT) followed by Agentic Reinforcement Learning with a novel gated reward function—to teach the model how to reason and when to use tools effectively.
Our main contributions are summarized as follows:
@@
  • A New Training Dataset and Evaluation Benchmark. We introduce a pipeline to curate, unify, and distill data from 10 existing sources, covering 12 different sports, including American football, badminton, baseball, basketball, boxing, diving, fencing, ice hockey, soccer, table tennis, volleyball, and gymnastics. In total, we construct 78K training data and 6.7K testing data.
  • Tool-Based Reinforcement Learning Strategy. We design a specialized reinforcement learning framework using Group Relative Policy Optimization (GRPO). We introduce a gated tool-use reward function that incentivizes the model to perform active, multi-turn visual reasoning while maintaining logical consistency.
  • State-of-the-Art Performance. Our DeepSport achieves an average score of 38.29, significantly outperforming all baselines, including general state-of-the-art (SOTA) models, and establishes a new SOTA for sports video tasks.
@@ 2 Related Work
In this domain, early sports video analysis relied on task-specific, traditional computer vision (CV) pipelines. For example, for player detection and tracking [cioppa2021cameracalibrationplayerlocalization, liu2023automatedplayeridentificationindexing, cioppa2022soccernet], rally prediction [xia2022vren, 10411638], automated commentary generation[huang2020generating, 10871604, 10.1145/3341105.3374063], fine-grained action recognition in specific sports [shao2020finegym, wu2022survey, WANG2023e18124, giancola2018soccernet]. While effective for their spesific tasks, these specialized models lack unified reasoning capabilities and cross-modal flexibility.
@@ 2.1 Video Understanding Using Multimodal Large Language Models
Early efforts in general video understanding extended MLLMs to handle video inputs, enabling video conversation and description (e.g., Video-ChatGPT [Maaz2023VideoChatGPT], Video-LLaVA [lin2023video]). However, these open source models usually fall behind SOTA commercial MLLMs, such as GPT [gpt4o], Gemini[geminiteam2024gemini], and Claude [Claude3.5] in most video tasks, especially those that require complicated reasoning on multiple video question answer benchmarks [fang2024mmbenchvideolongformmultishotbenchmark, zhou2025mlvubenchmarkingmultitasklong]. Recent developments have focused on combining reinforcement learning (RL) with MLLMs, pushing models to achieve stronger reasoning abilities and demonstrate enhanced performance on various video understanding tasks, such as Video-R1[feng2025video] and DeepVideo-R1 [park2025deepvideor1videoreinforcementfinetuning]. Inspired by the concept of “Thinking with images” proposed by OpenAI [a2025_thinking], recent works have begun using agentic reinforcement learning to train MLLM as an agent that can use tools to dynamically and iteratively interrogate visual content as part of its reasoning process [zhang2025landscapeagenticreinforcementlearning, fan2025grit, zheng2025deepeyesincentivizingthinkingimages]. This agentic concept has been extended to the temporal domain to handle long videos. Recent works introduce MLLMs that can “think with long videos” [zhang2025thinkingvideosmultimodaltoolaugmented, he2025framethinkerlearningthinklong]. These models act as agents that can densely sample new video frames on demand or perform multi-turn frame spotlighting by invoking a visual toolbox. This allows the model to actively “re-watch” segments and dynamically query the video for relevant frames, drastically reducing the number of frames processed and mitigating hallucination.
While these general-domain video MLLMs demonstrate powerful reasoning on benchmarks, such as Video-Holmes [cheng2025videoholmesmllmthinklike], they are trained for general-purpose rather than domain-specific tasks. The current general-domain MLLMs have been shown lacking the specialized domain knowledge required to understand the intricate rules, fine-grained actions, and professional terminology of diverse sports [xia2024sportu, rao2025multi].
@@ 2.2 Multimodal Large Language Models in Sports
Existing MLLMs in sports are typically limited in one of three ways: they are (i) training-free, relying on static knowledge graphs with limited tasks, such as no commentary generation task, which is one of the important things for sports viewers [chen2025finequestadaptiveknowledgeassistedsports]; (ii) single-sport, focusing on individual sports [rao2025multi, you2025timesoccer, kodathala2025sv33bsportsvideounderstanding, bao2025tennistvmultimodallargelanguage]; or (iii) single-task, such as models designed only for commentary [you2025timesoccer].
@@
Figure 2: DeepSport training overview. Given sport videos, we first perform data distillation with a teacher MLLM to construct DeepSport-CoT data and Supervised Fine-Tune a tool-augmented student model. We then further optimize the model with GRPO-based agentic reinforcement learning, where the agent iteratively calls a frame-extraction tool, produces chain-of-thought reasoning over new frames, and is guided by a reward manager that combines semantic accuracy, behavioral shaping, and format gating.
@@ 3 The DeepSport Framework
To address the gap in multi-task, multi-sport video understanding, we introduce DeepSport, an end-to-end trained MLLM trained with agentic reinforcement learning with a specific tool function, that extracts related frames and performs multi-turn generation. Our framework is built upon a two-stage training pipeline: (1) a Supervised Fine-Tuning (SFT) “cold-start” phase, followed by (2) an Agentic Reinforcement Learning (RL) phase. This SFT+RL paradigm is a validated approach for training capable “Thinking with Video” agents, as seen in recent works like FrameThinker [he2025framethinkerlearningthinklong]. Our methodology makes two primary contributions. First, we propose a data distillation pipeline to create a high-quality, large-scale dataset that includes the complex Chain-of-Thought (CoT) reasoning and tool-call syntax necessary for training the model. Second, we introduce a novel tool-based reward function within the Group Relative Policy Optimization (GRPO) framework to specifically incentivize effective multi-turn reasoning.
@@ 3.1 Deepsport Video Reasoning Paradigm
Our framework empowers the MLLM to “think with videos” [zhang2025thinkingvideosmultimodaltoolaugmented, he2025framethinkerlearningthinklong]. The framework engages in a multi-turn reasoning loop, actively and strategically seeking information from the video for its decision-making.
@@
  • Thought ($T_{t}$): The model’s textual reasoning, generated within <think>...</think> tags. The model generates this thought based on $F_{t}$ and all prior history.
@@
Inspired by the Cognitive Consistency Verification module in FrameThinker [he2025framethinkerlearningthinklong], we apply the following method to prevent of redundant exploration. If the model executes an action $A_{t}$ with the exact same temporal interval as its preceding action $A_{t-1}$, the framework treats this as an invalid “format error,” which is penalized during training.
@@ 3.2 Data Distillation Pipeline
1. Data Generation. We first curated a diverse set of 9 existing works (Table 1) with 10 different data sources to cover a wide spectrum of 12 distinct sports, spanning team sports (Soccer, Basketball, Volleyball, American Football, Ice Hockey, Baseball), racket games (Table Tennis, Badminton), combat sports (Fencing, Boxing), and artistic disciplines (Diving, Gymnastics). We unified these disparate data sources to establish a comprehensive taxonomy of tasks across four high-level dimensions: (1) Fine-Grained Recognition: encompassing foundational visual perception tasks such as Action Sequencing, Player Identification, Camera Transition Analysis, etc.; (2) Rule & Procedural Logic: focusing on regulatory knowledge, including Foul Classification, Referee Decision Analysis, Penalty Description, etc.; (3) Assessment & Coaching: requiring expert-level reasoning for Tactical Analysis, Technical Error Identification, Scoring Prediction, etc.; and (4) Live Commentary & Reporting: covering multimodal narration tasks like Play-by-Play Commentary, Game Situation Analysis, Score Reading, etc.
Since several of these benchmarks are not in a question-answering format (FineDiving [xu2022finediving] for action quality, SoccerReplay-1988 [rao2024towards] for commentary generation, FACTS [lai2024facts] for action classification, or T3Set [10.1145/3711896.3737407] for tactic correction), our first step was to unify them. We designed a series of task-specific templates to convert their original labels into structured QA pairs.
@@
SoccerBench [rao2025multi] Soccer Multi-task QA
SoccerReplay-1988 ([rao2024towards]) Soccer Commentary Generation (*)
SPORTU [xia2024sportu] Multi-sport Multi-level & -task QA
Sports-QA [li2024sports] Multi-sport Multi-task QA
SportR [xia2025sportrbenchmarkmultimodallarge] Multi-sport Foul & Tactic Reasoning QA
FineDiving [xu2022finediving] Diving Action Quality Assessment (*)
FACTS [lai2024facts] Fencing, Boxing Action Classification (*)
T3-Set ([10.1145/3711896.3737407]) Table Tennis Tactic Correction (*)
X-VARS([held2024x]) Soccer Foul Recognition & Explanation
2. Data Splitting and Integrity. We follow video-level splitting: all questions associated with a single video (even if there are multiple) are assigned exclusively to one data split. This prevents the model from being trained and tested on frames from the same video. Furthermore, we explicitly handle dataset overlap; for instance, any video clip used in the SoccerBench [rao2025multi] test set is excluded from our SoccerReplay-1988 [rao2024towards] training pool. The detailed number of QA split will be illustrated in Section 4.1.
3. CoT Data Distillation. With a unified set of QA prompts, we generate high-quality reasoning trajectories. We prompt the current SOTA model, Qwen3-VL-235B-A22B-Thinking [qwen3technicalreport], to generate detailed, step-by-step reasoning. Critically, the prompt instructs the model to invoke the frame_extraction_tool (defined in Section 3.1) when necessary. This process results in CoT trajectories that are interleaved with tool calls.
We employ an LLM-as-Judge (using DeepSeek-V3.2-Exp [deepseekai2024deepseekv32]) to filter for high-quality trajectories. The judge scores each generated CoT on the predicted answer. Only trajectories that pass a high-quality threshold are retained for SFT Training. As a result, we constructed the DeepSport-CoT-15k dataset, which contains 15,000 high-quality Q&A pairs with CoT annotations, to be used later in SFT training.

### v2

@@ Abstract
Sports video understanding requires perceiving high-speed dynamics, complex rules, and long temporal contexts. Yet, current Multimodal Large Language Models (MLLMs) remain narrowly focused on single sports, specific tasks, or training-free paradigms. We introduce DeepSport, the first end-to-end trained MLLM for multi-task, multi-sport video understanding. DeepSport shifts from passive frame processing to active, iterative reasoning, dynamically extracting frames to “think with videos.” To train our model, we curate a unified 78k-sample dataset via a rigorous three-step text-and-vision distillation pipeline. We then employ a progressive two-stage training strategy: a Sports Curriculum Supervised Fine-Tuning phase to build foundational perception, followed by Agentic Reinforcement Learning with a novel tool-use reward. Extensive experiments on a comprehensive 6.7k benchmark demonstrate that DeepSport achieves state-of-the-art performance—outperforming powerful proprietary and open-source models—while utilizing significantly fewer frames. Furthermore, it exhibits strong zero-shot transferability to unseen sports and broad motion recognition tasks, establishing a highly efficient and generalized foundation for complex video reasoning.
11footnotetext: * Equal contribution.
@@ 1 Introduction
The domain of sports analysis has long served as a fertile ground for interdisciplinary research, driving advancements in both computer vision and natural language processing [51]. Traditional vision approaches have achieved remarkable success in specific, isolated tasks, such as fine-grained action detection and recognition [46, 20], event spotting [60], highlight generation [36, 37], and player and ball trajectory detection and tracking in team sports [8, 52, 63, 16]. Simultaneously, NLP techniques have been effectively deployed for automated match reporting [25] and sentiment analysis [53]. Recently, the emergence of Multimodal Large Language Models (MLLMs) [38, 19] has accelerated this trend, enabling more complex applications in sports [65], such as AI-assisted refereeing [22], trajectory generation [5], and commentary generation [44, 70].
In response to these capabilities, the community has produced a surge of high-quality, public benchmarks covering a wide array of disciplines. These range from action quality assessment in diving [67] and fine-grained player skill estimation in basketball [40], to tactical classification in fencing [28], and comprehensive question answering in large-scale benchmarks [64, 66, 42]. These datasets provide the necessary foundation for evaluating sophisticated sports intelligence.
Despite the creation of data and benchmarks, the development of capable models has lagged behind. Current sports-specific MLLMs exhibit a significant fragmentation. For example, trained models are predominantly soccer-centric [42, 70], optimizing for specific tasks such as commentary generation within a single sport, but failing to generalize to others. On the other hand, attempts at multi-sport reasoning, such as FineQuest [6], rely on training-free paradigms. Consequently, to the best of our knowledge, a unified, end-to-end trained MLLM capable of performing multiple tasks (e.g., commentary, foul detection, QA) across diverse sports does not yet exist.
To bridge this gap, we introduce DeepSport, the first sports-specific MLLM framework explicitly trained for multi-task, multi-sport video understanding (Figure 1). Inspired by the emerging paradigm of “Thinking with Videos” [72, 21, 18], DeepSport shifts from passive frame processing to active, iterative video interrogation via a specialized frame-extraction tool. To enable this, we devise a rigorous data distillation pipeline that synthesizes high-quality Chain-of-Thought (CoT) trajectories from 10 diverse data sources covering 12 different sports. We then employ a progressive two-stage strategy—a Sports Curriculum Supervised Fine-Tuning (SFT), inspired by the nature of sports, to incrementally build foundational perception, followed by Agentic Reinforcement Learning with a novel gated reward function—to teach the model how to reason and when to use tools effectively. Our main contributions are summarized as follows:
@@
  • A New Benchmark and Rigorous Distillation Pipeline. We introduce a pipeline to curate, unify, and distill data from 10 existing sources across 12 sports. By utilizing a stringent three-step LLM/VLM filtering mechanism to guarantee logic and retrieval consistency, we construct 78K high-quality training trajectories and a testing benchmark with 6.7K samples.
  • Curriculum SFT and Tool-Based RL Strategy. Inspired by the nature and specificity of sports, we design a progressive curriculum-learning strategy during SFT to build visual grounding before high-level logic. This is followed by a specialized GRPO reinforcement learning framework with a gated tool-use reward function that incentivizes active, multi-turn visual reasoning.
  • SOTA Performance and Broad Generalizability. DeepSport achieves an average score of 37.67, establishing a new SOTA for sports video tasks. Furthermore, it demonstrates robust zero-shot transferability to unseen sports and broad motion recognition tasks using significantly fewer frames, proving our paradigm internalizes underlying athletic mechanics rather than merely overfitting to specific rules.
@@ 2 Related Work
In this domain, early sports video analysis relied on task-specific, traditional computer vision (CV) pipelines. For example, for player detection and tracking [8, 31, 9], rally prediction [62, 63], automated commentary generation[25, 68, 26], fine-grained action recognition in specific sports [46, 58, 54, 20]. While effective for their specific tasks, these specialized models lack unified reasoning capabilities and cross-modal flexibility.
@@ 2.1 Video Understanding Using Multimodal Large Language Models
Early efforts in general video understanding extended MLLMs to handle video inputs, enabling video conversation and description (e.g., Video-ChatGPT [35], Video-LLaVA [30]). However, these open source models usually fall behind SOTA commercial MLLMs, such as GPT [38], Gemini[19], and Claude [1] in most video tasks, especially those that require complicated reasoning on multiple video question answer benchmarks [13, 74]. Recent developments have focused on combining reinforcement learning (RL) with MLLMs, pushing models to achieve stronger reasoning abilities and demonstrate enhanced performance on various video understanding tasks, such as Video-R1[14] and DeepVideo-R1 [41]. Inspired by the concept of “Thinking with images” proposed by OpenAI [39], recent works have begun using agentic reinforcement learning to train MLLM as an agent that can use tools to dynamically and iteratively interrogate visual content as part of its reasoning process [71, 12, 73]. This agentic concept has been extended to the temporal domain to handle long videos. Recent works introduce MLLMs that can “think with long videos” [72, 21]. These models act as agents that can densely sample new video frames on demand or perform multi-turn frame spotlighting by invoking a visual toolbox. This allows the model to actively “re-watch” segments and dynamically query the video for relevant frames, drastically reducing the number of frames processed and mitigating hallucination.
While these general-domain video MLLMs demonstrate powerful reasoning on benchmarks, such as Video-Holmes [7], they are trained for general-purpose rather than domain-specific tasks. The current general-domain MLLMs have been shown lacking the specialized domain knowledge required to understand the intricate rules, fine-grained actions, and professional terminology of diverse sports [66, 42].
@@ 2.2 Multimodal Large Language Models in Sports
Existing MLLMs in sports are typically limited in one of three ways: they are (i) training-free, relying on static knowledge graphs with limited tasks, such as no commentary generation task, which is one of the important things for sports viewers [6]; (ii) single-sport, focusing on individual sports [42, 70, 27, 4]; or (iii) single-task, such as models designed only for commentary [70].
@@
Figure 2: DeepSport training overview. Given sports videos, we first perform data distillation with a teacher MLLM to construct DeepSport-CoT data and Curriculum Supervised Fine-Tune model. We then further optimize the model with GRPO-based agentic reinforcement learning, where the agent iteratively calls a frame-extraction tool, produces chain-of-thought reasoning over new frames, and is guided by a reward manager that combines semantic accuracy, behavioral shaping, and format gating.
@@ 3 The DeepSport Framework
To address the gap in multi-task, multi-sport video understanding, we introduce DeepSport, an end-to-end trained MLLM trained with agentic reinforcement learning with a specific tool function, that extracts related frames and performs multi-turn generation. Our framework is built upon a two-stage training pipeline: (1) a Supervised Fine-Tuning (SFT) “cold-start” phase, followed by (2) an Agentic Reinforcement Learning (RL) phase. This SFT+RL paradigm is a validated approach for training capable “Thinking with Video” agents, as seen in recent works like FrameThinker [21]. Our methodology makes two primary contributions. First, we propose a data distillation pipeline to create a high-quality, large-scale dataset that includes the complex Chain-of-Thought (CoT) reasoning and tool-call syntax necessary for training the model. Second, we introduce a novel tool-based reward function within the Group Relative Policy Optimization (GRPO) framework to specifically incentivize effective multi-turn reasoning.
@@ 3.1 Deepsport Video Reasoning Paradigm
Our framework empowers the MLLM to “think with videos” [72, 21]. The framework engages in a multi-turn reasoning loop, actively and strategically seeking information from the video for its decision-making.
@@
  • Thought ($T_{t}$): The model’s textual reasoning, generated within
    <think>...</think> tags. The model generates this thought based on $F_{t}$ and all prior history.
@@
Inspired by the Cognitive Consistency Verification module in FrameThinker [21], we apply the following method to prevent of redundant exploration. If the model executes an action $A_{t}$ with the exact same temporal interval as its preceding action $A_{t-1}$, the framework treats this as an invalid “format error,” which is penalized during training.
@@ 3.2 Data Distillation Pipeline
1. Data Generation. We first curated a diverse set of 9 existing works (Table 1) with 10 different data sources to cover a wide spectrum of 12 distinct sports, spanning team sports (Soccer, Basketball, Volleyball, American Football, Ice Hockey, Baseball), racket games (Table Tennis, Badminton), combat sports (Fencing, Boxing), and artistic disciplines (Diving, Gymnastics). We unified these disparate data to establish a comprehensive taxonomy of tasks across four high-level dimensions: (1) Fine-Grained Recognition: encompassing foundational visual perception tasks such as Action Sequencing, Player Identification, Camera Transition Analysis, etc.; (2) Rule & Procedural Logic: focusing on regulatory knowledge, including Foul Classification, Referee Decision Analysis, Penalty Description, etc.; (3) Assessment & Coaching: requiring expert-level reasoning for Tactical Analysis, Technical Error Identification, Scoring Prediction, etc.; (4) Live Commentary & Reporting: covering multimodal narration tasks like Play-by-Play Commentary, Game Situation Analysis, Score Reading, etc.
Since several of these benchmarks are not in a question-answering format (FineDiving [67] for action quality, SoccerReplay-1988 [43] for commentary generation, FACTS [28] for action classification, or T3Set [34] for tactic correction), our first step was to unify them. We designed a series of task-specific templates to convert their original labels into structured QA pairs.
@@
SoccerBench [42] Soccer Multi-task QA
SoccerReplay-1988 ([43]) Soccer Commentary Generation (*)
SPORTU [66] Multi-sport Multi-level & -task QA
Sports-QA [29] Multi-sport Multi-task QA
SportR [61] Multi-sport Foul & Tactic Reasoning QA
FineDiving [67] Diving Action Quality Assessment (*)
FACTS [28] Fencing, Boxing Action Classification (*)
T3-Set ([34]) Table Tennis Tactic Correction (*)
X-VARS([23]) Soccer Foul Recognition & Explanation
2. Data Splitting and Integrity. We follow video-level splitting: all questions associated with a single video (even if there are multiple) are assigned exclusively to one data split. This prevents the model from being trained and tested on frames from the same video. Furthermore, we explicitly handle dataset overlap; for instance, any video clip used in the SoccerBench [42] test set is excluded from our SoccerReplay-1988 [43] training pool. The detailed number of QA split will be illustrated in Section 4.1.
3. CoT Data Distillation. With a unified set of QA prompts, we generate high-quality reasoning trajectories. We prompt the current SOTA model, Qwen3-VL-235B-A22B-Thinking [50], to generate detailed, step-by-step reasoning. Critically, the prompt instructs the model to invoke the frame_extraction_tool (defined in Section 3.1) when necessary. This process results in CoT trajectories that are interleaved with tool calls.
To ensure the data quality and prevent reasoning hallucinations or redundant tool usage, we propose a rigorous three-step filtering pipeline: Step 1 - Accuracy Check: We use DeepSeek-V3.2-Exp [11] to verify if the model’s predicted final answer matches the ground truth. Step 2 - Logic Consistency Review: We employ DeepSeek-V3.2 [11] to evaluate the reasoning trace. This ensures the reasoning is self-consistent and that any revision of previous conclusions is justified by new information, strictly penalizing unprompted contradictions. Step 3 - Retrieval Usefulness Evaluation: We utilize a Kimi K2.5 [49] to assess the visual necessity of the tool call. It verifies whether the dynamically retrieved frames provide genuinely new, finer-grained visual details compared to the initial frames, ensuring the retrieval is not redundant.
Only trajectories that successfully pass all three verification steps are retained. As a result, we constructed the DeepSport-CoT-14K dataset, which contains 14,599 high-quality Q&A pairs with robust CoT annotations, to be used later in SFT training.